# HARBOR: Holistic Adaptive Risk assessment model for BehaviORal healthcare

### v2

@@ 1 Introduction
Accurate assessment of mental health risk is foundational to effective psychiatric and therapeutic care. Clinicians routinely integrate heterogeneous signals—sleep, activity, metabolic health, self-reported questionnaires, and lived context—into qualitative judgments about patient mood and risk. Automating or augmenting this process remains difficult, particularly when predictions must be discrete, interpretable, and temporally grounded.
@@
Recent advances in large language models (LLMs) suggest promise in reasoning over structured and semi-structured health data. However, most prior work evaluates LLMs on open-ended clinical question answering rather than calibrated risk scoring. Furthermore, little work examines whether general-purpose LLMs can reliably predict longitudinal mood trajectories from compact behavioral feature sets.
@@
2 PEARL
  • Body composition: weight, body fat percentage
  • Behavioral proxies: number of photos taken, location entropy
  • Financial context: monthly expenses normalized by income
  • Clinical questionnaires: PHQ-9 and GAD-7
@@ 3.3 Self-Taught Reasoning
To improve structured reasoning over tabular inputs, we employ a self-taught reasoning (STaR) approach, where the model iteratively generates and refines its own reasoning traces during training.
@@ 6 Calibration
7 Interpretability
@@ 10 Related Work
The challenge of predicting mental health outcomes has long been recognized in psychiatry. Classical work by Meehl demonstrated that simple statistical models can outperform clinical judgment in behavioral prediction, a result that has replicated across decades of clinical domains (meehl1954clinical). More recently, large-scale meta-analyses have shown that traditional psychiatric risk factors—particularly for suicide—have limited predictive power, motivating the use of machine learning–based risk models (franklin2017risk). In contrast to unstructured clinical judgment, structured risk scores such as the National Early Warning Score 2 (NEWS2) have seen widespread adoption in general medicine by aggregating physiological signals into an interpretable, discrete score for clinical decision-making (royal2017news2). However, comparable standardized scoring systems for behavioral health remain limited.
@@
Recent work has explored predictive modeling using structured electronic health records (EHRs), demonstrating improved performance for outcomes such as psychiatric readmission and suicide attempts (simon2018suicide; kessler2020predicting). Beyond EHRs, advances in mobile sensing and digital phenotyping have enabled continuous, longitudinal measurement of behavioral signals such as sleep, activity, mobility, and self-reported mood (onella2016mobile). Publicly released datasets capturing such signals have supported modeling of mood dynamics and relapse risk in real-world settings (pratap2020all; torous2021digital). These studies highlight the importance of combining physiological, behavioral, and self-report features—an approach directly reflected in the PEARL dataset.
@@
Large language models (LLMs) have recently been explored for clinical applications, including medical question answering, summarization, and decision support (singhal2023large). Early studies suggest that LLMs can perform competitively on medical reasoning benchmarks, yet their reliability for calibrated risk prediction remains unclear (nori2023capabilities). In psychiatry, LLMs have been proposed for tasks such as mental health screening, therapy assistance, and patient engagement, but existing evaluations remain limited in scale and standardization (bickmore2018automated; torous2023trustworthy). Importantly, prior work largely focuses on free-text interaction rather than discrete, interpretable risk scoring.
10.1 Advances in Large Language Models
Large language models (LLMs) have undergone rapid advances over the past decade, driven by scaling laws, architectural innovations, and improved training paradigms. Early neural language models based on recurrent architectures demonstrated the feasibility of learning distributed representations of language, but were limited in their ability to model long-range dependencies (mikolov2010recurrent). The introduction of the Transformer architecture fundamentally altered this trajectory by enabling efficient attention-based modeling of global context, forming the foundation for modern LLMs (vaswani2017attention).
Subsequent work demonstrated that increasing model and dataset scale leads to emergent capabilities in reasoning, abstraction, and generalization (kaplan2020scaling; hoffmann2022training). Large pretrained models such as GPT-3 showed that a single model could perform a wide variety of tasks through prompting alone, without task-specific fine-tuning (brown2020language). This observation led to extensive research on in-context learning, instruction tuning, and alignment, substantially improving usability and controllability (wei2022finetuned; ouyang2022training).
More recently, LLMs have achieved state-of-the-art performance across a range of professional and academic benchmarks, including medicine. Models such as PaLM and GPT-4 demonstrate strong performance on medical question answering and clinical reasoning tasks, approaching or exceeding expert-level accuracy in constrained settings (openai2023gpt4). These results suggest that LLMs encode substantial domain knowledge, even when trained primarily on general-purpose corpora.
However, several limitations remain salient for clinical deployment. First, LLMs are known to be poorly calibrated, often expressing high confidence in incorrect outputs (guo2017calibration). Second, most evaluations focus on free-text generation or multiple-choice reasoning rather than discrete, clinically interpretable risk scores. Third, general-purpose LLMs lack explicit alignment with functional impairment frameworks commonly used in psychiatry, limiting their utility in real-world decision support (torous2023trustworthy).
Recent work has begun exploring domain-adapted and instruction-specialized models for healthcare, emphasizing safety, interpretability, and controlled outputs (singhal2023large). Nonetheless, existing systems largely operate as black-box assistants, providing recommendations without transparent mappings to clinical constructs such as symptom severity or work impairment.
In contrast, HARBOR builds on advances in large-scale language modeling while explicitly constraining outputs to a clinically grounded, discrete risk scale. By combining domain-specific training, calibrated confidence estimation, and interpretable score definitions, HARBOR addresses key gaps that remain unresolved by general-purpose LLMs in behavioral healthcare settings.
Positioning of This Work.
HARBOR builds on these lines of research by combining structured behavioral health features with domain-adapted language modeling to produce a calibrated, discrete mood risk score. Unlike prior LLM-based systems, HARBOR is explicitly evaluated against traditional machine learning baselines and standardized across multiple ablation dimensions, positioning it as a step toward interpretable and reproducible behavioral health risk assessment.
@@ 11 Conclusion and Future Work
Future work includes expanding PEARL to more patients, increasing temporal resolution to daily or hourly predictions, incorporating passive sensor data, and exploring HARBOR as a training and decision-support tool for clinicians and mental health professionals.

### v3

@@ 1 Introduction
Accurate assessment of mental health risk is foundational to effective psychiatric and therapeutic care. Clinicians routinely integrate heterogeneous signals—sleep, activity, metabolic health, self-reported questionnaires, and lived context—into qualitative judgments about patient mood and risk. Automating or augmenting this process remains difficult, particularly when predictions must be discrete, interpretable, and temporally grounded. Recent advances in large language models (LLMs) suggest promise in reasoning over structured and semi-structured health data. However, most prior work evaluates LLMs on open-ended clinical question answering rather than calibrated risk scoring. Furthermore, little work examines whether general-purpose LLMs can reliably predict longitudinal mood trajectories from compact behavioral feature sets.
@@ 2.3 Self-Taught Reasoning
To improve structured reasoning over tabular inputs, we employ a self-taught reasoning (STaR) approach, where the model iteratively generates and refines its own reasoning traces during training (zelikman2022star).
3 PEARL
  • Body composition: weight, body fat percentage
  • Behavioral proxies: number of photos taken, location entropy
  • Financial context: monthly expenses normalized by income
  • Clinical questionnaires: PHQ-9 and GAD-7
3.1 Ethical Considerations
3.2 Dataset Splits
@@ 7 Calibration
8 Imputation
Although PEARL is largely dense, real-world behavioral health data is often missing or intermittently observed. To assess robustness under missingness, we simulate sparsity by masking a subset of feature values at random and then imputing them prior to inference. We compare four imputation strategies spanning classical statistical baselines, iterative multivariate methods, and model-based generation.
  • Median/Mode Imputation: Replaces missing numeric values with the training-set median and missing categorical values with the most frequent category.
  • Regression Imputation: Predicts each missing feature using a regression model fit on observed features, then fills missing values with the model’s predictions.
  • MICE: Uses Multiple Imputation by Chained Equations, iteratively imputing each feature conditional on the others and averaging across multiple imputations.
  • LLM-Generated Imputation: Prompts an LLM to generate plausible missing feature values conditioned on the observed fields and basic clinical plausibility constraints.
Table 8 reports accuracy under increasing missingness rates. Across all masking levels, MICE performs best, followed by regression imputation, then median/mode. LLM-generated imputations perform worst, though the gap is modest, suggesting that constrained generation can be a viable fallback when classical assumptions fail.
Table 8: Accuracy under Simulated Missingness with Different Imputation Methods (Simulated).
Imputation Method 10% Missing 25% Missing 40% Missing
Median/Mode 0.67 0.62 0.57
Regression 0.68 0.64 0.59
MICE 0.70 0.66 0.61
LLM-Generated 0.66 0.61 0.56
@@ 11 Related Work
The challenge of predicting mental health outcomes has long been recognized in psychiatry. Classical work by Meehl demonstrated that simple statistical models can outperform clinical judgment in behavioral prediction, a result that has replicated across decades of clinical domains (meehl1954clinical). More recently, large-scale meta-analyses have shown that traditional psychiatric risk factors—particularly for suicide—have limited predictive power, motivating the use of machine learning–based risk models (franklin2017risk). In contrast to unstructured clinical judgment, structured risk scores such as the National Early Warning Score 2 (NEWS2) have seen widespread adoption in general medicine by aggregating physiological signals into an interpretable, discrete score for clinical decision-making (smith2019national). However, comparable standardized scoring systems for behavioral health remain limited.
Subsequent work has highlighted the inherent difficulty of psychiatric risk prediction, particularly for outcomes such as suicide attempts, relapse, or mood destabilization. Large cohort studies and systematic reviews consistently report low positive predictive value for individual risk factors, even when statistically significant, underscoring the need for multivariate and longitudinal modeling approaches (franklin2017risk; kessler2020predicting). As a result, recent research has shifted from single-factor screening toward composite risk scores that integrate behavioral, physiological, and contextual signals. In parallel, concerns have been raised regarding the interpretability and clinical acceptability of black-box risk models. Studies show that clinicians are more likely to trust and adopt decision-support tools that expose clinically meaningful intermediate representations, such as discrete risk categories or functional impairment levels, rather than continuous opaque scores (caruana2015intelligible). These findings motivate the design of structured, interpretable risk scales such as the Harbor Risk Score.
@@
Recent work has explored predictive modeling using structured electronic health records (EHRs), demonstrating improved performance for outcomes such as psychiatric readmission and suicide attempts (simon2018suicide; kessler2020predicting). Beyond EHRs, advances in mobile sensing and digital phenotyping have enabled continuous, longitudinal measurement of behavioral signals such as sleep, activity, mobility, and self-reported mood (felix2019mobile). Publicly released datasets capturing such signals have supported modeling of mood dynamics and relapse risk in real-world settings (pratap2019accuracy; melcher2020digital). Recent advances in digital phenotyping have enabled continuous collection of behavioral signals via smartphones and wearables, including sleep, activity, mobility, and social interaction proxies (onella2016mobile). These studies highlight the importance of combining physiological, behavioral, and self-report features—an approach directly reflected in the PEARL dataset. However, many existing datasets are either short-term, sparsely labeled, or lack clinician-validated ground truth, limiting their utility for model development and evaluation.
Advances in Large Language Models
Since 2022, progress in large language models (LLMs) has been driven by improved training recipes, instruction tuning, and alignment, alongside continued gains from scaling under compute-optimal regimes (hoffmann2022training; wei2022finetuned; ouyang2022training). Foundational demonstrations such as GPT-3 established the viability of broad task competence via prompting (brown2020language), while newer frontier systems have expanded capabilities in multimodal reasoning and long-context retrieval—most notably Gemini and Gemini 1.5, which report effective reasoning over very long contexts and improved performance on long-document tasks (geminiteam2023gemini; geminiteam2024gemini15). In medicine, recent evaluations show strong performance of instruction-following LLMs on constrained clinical reasoning and question answering benchmarks, motivating their use in decision-support settings (openai2023gpt4; singhal2023large). However, important limitations remain salient for deployment: language model probability outputs can be poorly calibrated even in controlled settings (lovering2024lmcalibration), and most clinical evaluations still emphasize free-text responses rather than discrete, clinically interpretable risk scores aligned with functional impairment and workflow constraints (torous2025assessing). HARBOR builds on these advances while explicitly constraining outputs to a clinically grounded discrete risk scale with confidence estimation, targeting the gap between general capabilities and deployable behavioral health risk stratification.
@@
Large language models (LLMs) have recently been explored for clinical applications, including medical question answering, summarization, and decision support (singhal2023large). Early studies suggest that LLMs can perform competitively on medical reasoning benchmarks, yet their reliability for calibrated risk prediction remains unclear (nori2023capabilities). In psychiatry, LLMs have been proposed for tasks such as mental health screening, therapy assistance, and patient engagement, but existing evaluations remain limited in scale and standardization (bickmore2018automated; torous2025assessing). Importantly, prior work largely focuses on free-text interaction rather than discrete, interpretable risk scoring.
12 Positioning of this Paper
Position: Behavioral health machine learning should prioritize clinically grounded, calibrated, and interpretable discrete risk scoring, and should recognize small, deeply validated longitudinal datasets as essential scaffolding for responsible evaluation and deployment.
Behavioral healthcare presents a uniquely high-stakes setting for ML, where predictions must be interpretable, uncertainty-aware, and aligned with clinical workflow. In such settings, raw accuracy improvements on large, weakly labeled datasets are often less actionable than systems that produce calibrated, discrete risk categories tied to functional impairment. We argue that progress in this domain requires shifting emphasis away from unconstrained free-text outputs toward structured decision-support primitives that clinicians can reliably interpret, audit, and act upon.
Accordingly, this work is intentionally framed as a position paper and proof-of-concept rather than a definitive empirical study. The PEARL dataset is small (three patients) but deeply curated and provider-validated, and is not intended to support population-level generalization claims. Instead, it serves to illustrate a core methodological position: meaningful progress in behavioral health ML is bottlenecked less by model capacity than by the absence of clinically validated longitudinal datasets and standardized evaluation protocols for calibrated risk scoring.
12.1 Alternative Views
A credible alternative view is that research effort should focus primarily on large-scale datasets, with calibration and interpretability addressed only after sufficient statistical power is achieved. While we agree that scale is ultimately necessary, we argue that scaling without a well-defined, clinically grounded target risks optimizing metrics that do not translate to real-world decision-making. Another alternative view favors free-text clinical assistants over discrete risk scores. We counter that discrete, calibrated outputs remain central to triage, escalation, and accountability in clinical practice, and provide clearer affordances for safety and monitoring.
12.2 Call to Action
We call on the community to: (i) invest in longitudinal behavioral health datasets designed explicitly for discrete risk scoring and calibration analysis; (ii) adopt evaluation standards that emphasize calibration, robustness to missingness, and temporal and patient-level generalization; and (iii) prioritize constrained, interpretable model outputs suitable for clinical decision support. Under this framing, our contribution is to advocat
@@ 13 Conclusion and Future Work
Future work includes expanding PEARL to more patients, increasing temporal resolution to daily or hourly predictions, and exploring HARBOR as a training and decision-support tool for clinicians and mental health professionals.